OpenStudy (anonymous):

Find the value of b so that vectors U = 5i+12j and V = 4i+bj are perpendicular

OpenStudy (turingtest):

when two vectors \(\vec v\) and \(\vec u\) are perpendicular, their dot product is zero

OpenStudy (anonymous):

okk

OpenStudy (turingtest):

hence just take the dot product and set it to zero\[\vec u\cdot\vec v=0\]and solve for \(b\)

OpenStudy (anonymous):

Would the Dot product be 32 ?

OpenStudy (anonymous):

I took 5x4 and 12x1 i don't know

OpenStudy (turingtest):

the dot product can't be non-zero or else the two vectors would not be perpendicular, as I mentioned above. why did you do 12x1 ? the j-component of the second vector is our unknown, \(b\), not 1 try again keeping the variable \(b\) where it belongs

OpenStudy (anonymous):

i got 20+12bj^2 is that ok ?

OpenStudy (turingtest):

yes, but\[(\hat j)^2=\hat j\cdot\hat j=1\]so you can simplify that a bit and what are we setting that expression equal to again?

OpenStudy (anonymous):

b ?

OpenStudy (turingtest):

no, two vectors are perpendicular if their dot product is...?

OpenStudy (anonymous):

0

OpenStudy (turingtest):

yes, so what are we solving? write the whole equation please

OpenStudy (anonymous):

ohhhhh i got it,

OpenStudy (turingtest):

:)

OpenStudy (anonymous):

0 = 20+12bj^2

OpenStudy (anonymous):

-20/12

OpenStudy (anonymous):

oh i see haha :))

OpenStudy (turingtest):

yup :D nice job! (again j^2=1)

OpenStudy (anonymous):

ohhh is it always like that ?

OpenStudy (turingtest):

is j^2 always 1 you mean?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

yes, because \(\hat j=\langle0,1\rangle\), so \((\hat j)^2=\hat j\cdot\hat j=\langle0,1\rangle\cdot\langle0,1\rangle=1\) (for the same reason, the \(\hat i\) disappeared)

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

thanks @TuringTest

OpenStudy (turingtest):

anytime @Ala123 :D